### v1

@@ 4 Application to Cryptocurrency Returns
Figure 3: Daily parameter estimation using NBE for 2022. Each row represents bitcoin (BTC), ethereum (EHT) and ripple (XRP) respectively. The left most panel represents the log returns data. The following panels are the estimates of $\sigma^{2}$, $\alpha_{1}$, and $\alpha_{2}$, of a level $2$ DVG process, respectively. The dashed lines corresponds to Jan 25, May 11, Jun 14, and Nov 11.
@@
Figure 4 provides further insight by comparing characteristic functions and probability density functions for the selected high-volatility days. These plots illustrate the flexibility of the NBE approach in adapting to changing market conditions. The day-to-day variation in the characteristic and density functions underscores the non-stationary nature of financial data, and how the model dynamically adjusts to capture this complexity.
@@
Figure 5 offers a visual representation of the $90\%$ confidence intervals that surround the parameter estimates. Notably, the non-overlapping confidence intervals throughout the year indicate that the underlying process is indeed non-stationary – its parameters evolve meaningfully over time.
Moreover, the estimated confidence intervals allow us to explore comparative behavior across cryptocurrencies. In particular, we define two cryptocurrencies as having different distributions on a given day if at least one of their corresponding confidence intervals does not overlap. Applying this criterion, we find that Bitcoin and XRP exhibit different distributions on 239 days in 2022, corresponding to roughly $65\%$ of the year. Similarly, Bitcoin and Ethereum differ on approximately $40\%$ of the days, and Ethereum and XRP differ on about $38\%$ of the days.
@@
Time [s] 45,552 656 5
Improvement (%) 99.99% 99.94% -
Speed-up (%) 910,940% 13,020% -
Table 4: Time needed to perform daily parameter estimation for the year 2022 averaged across 3 different cryptocurrencies. The second row contains the improvement time of NBE in comparison to the other two methods, and the third row shows the speed-up in estimation by using NBE.
@@
In addition to accuracy and robustness, computational efficiency is a key consideration for practical deployment. Table 4 compares the runtime of NBE against MELE and least squares. NBE is approximately 9,000 times faster than MELE and 130 times faster than least squares, making it an appealing option for real-time applications.
@@ 4.1 The Long-Term Evolution of Bitcoin
We also analyzed the long-term evolution of DVG process parameters fitted to BTC over a nine-year period spanning 2014 to 2022. The NBE approach shines in this application, with nearly a decade of data requiring only approximate 45 seconds to compute daily estimates for the entire period. Figure 6 displays these daily estimates for $\sigma^{2}$, $\alpha_{1}$, and $\alpha_{2}$ using the NBE framework. The parameters show clear signs of non-stationarity, with periods of relative stability punctuated by bursts of volatility. Since the variance estimates are difficult to interpret directly due to scale differences, we additionally visualize their evolution in log-scale in Figure 7.
@@ 5 Discussion
NBE is a flexible, amortized framework for parameter estimation that approximates the Bayes risk using a neural network trained via simulation. This allows NBE to bypass the need for closed-form likelihoods, making it especially well-suited for complex models such as Lévy processes, where traditional methods are usually limited by computational cost.
Our simulation studies showed that NBE consistently achieves lower estimation error across all parameters compared to established techniques like least squares and maximum empirical likelihood. Moreover, it does so with dramatically reduced computation times, achieving speedups of several orders of magnitude, making it a compelling option for large-scale or time-sensitive applications.
We further demonstrated the utility of NBE on high-frequency cryptocurrency return data. The method effectively captured temporal changes in process parameters, responded sensitively to market events, and offered robust uncertainty quantification. These findings highlight NBE’s capacity to adapt to the non-stationary nature of real-world financial data while delivering stable and interpretable estimates. Moreover, we were able to obtain parameter estimations for high-frequency 1-minute bitcoin data for 9 years in matter of seconds.
Nonetheless, certain limitations remain. Accurately estimating parameters associated with deep subordination, such as those in the level-2 DVG model, proved more difficult in low-signal regimes. Future work may focus on improving estimation for these challenging regimes, possibly by incorporating structured priors or attention mechanisms to enhance model sensitivity. Future directions may also include generalizing the method to multivariate Lévy processes, or exploring extrapolation beyond the prior distribution of the parameters.
@@ Acknowledgements
This research was supported by National Science Foundation grant DMS-2310487.
@@ 6.1 Simulation studies
To train the estimator across a broad range of plausible values, we selected reasonably uninformative prior distributions for the parameters of each process, similar as the priors used in 3.3. These were chosen to reflect realistic yet flexible ranges for each model’s parameter space.
  • Compound Poisson process:
    $\lambda\sim U(0.1,1.3),\qquad\mu\sim U(-0.6,0.6),\qquad\sigma^{2}\sim U(10^{-3},0.3)$
  • Merton jump-diffusion model:
    $\mu\sim U(-0.8,0.8),\,\sigma^{2}\sim U(10^{-3},1),\,\lambda\sim U(0.1,1.5),\,\mu_{J}\sim U(-1.5,1.5),\,\sigma_{J}^{2}\sim U(0.1,1.7)$
  • Variance gamma process:
    $\gamma\sim U(-1.5,1.5),\qquad\sigma^{2}\sim U(10^{-4},2),\qquad\alpha\sim U(0.1,3)$
  • Deep variance gamma (level 2):
    $\sigma^{2}\sim U(10^{-6},3),\qquad\alpha_{1},\alpha_{2}\sim U(10^{-6},25)$
@@ 6.1.1 Experiments on $n_{t}$
We began our investigation by fixing $K=5{,}000$ and $J=10$, and exploring the impact of $n_{t}$—the number of observations per dataset—on both training time and estimation accuracy. Specifically, we evaluated how varying $n_{t}$ affects the root mean squared error (RMSE) of the estimated parameters and the overall training time of the neural network.
To ensure robustness, we repeated this analysis across the three Lévy process models discussed in 3.2. For each model and each value of $n_{t}$, the estimator was trained on $K\cdot J$ datasets and then evaluated on an independent test set consisting of $1,000$ simulated datasets.
To visualize the results, we plotted RMSE against $n_{t}$ for each parameter in each model. These plots help identify the point at which increasing the input length yields little or no additional improvement in accuracy—what we refer to as a point of diminishing returns.
Across all models, we observe that RMSE steadily improves as $n_{t}$ increases, with diminishing returns beyond approximately 3,00030003{,}0003 , $0005{,}$ 000 observations. In other words, increasing $n_{t}$ beyond this range yields minimal gains in accuracy, as shown in Figure 8. The figure also presents results for two different models—the compound Poisson and variance gamma processes—demonstrating similar trends. In fact, for the variance gamma process (bottom row), the point of diminishing returns appears even earlier, around $1{,}500$ observations.
Given these findings, we selected $n_{t}=1{,}440$ in our final experiments. This value corresponds to the number of one-minute log returns in a single trading day (24 hours), making it a natural and interpretable choice for our cryptocurrency data application. While it lies slightly below the saturation threshold, it strikes a practical balance between empirical performance and application-specific relevance.

### v2

@@ 4 Application to Cryptocurrency Returns
Figure 3: Daily parameter estimation using NBE for 2022. Each row represents bitcoin (BTC), ethereum (EHT) and ripple (XRP) respectively. The left most panel represents the log returns data. The following panels are the estimates of $\sigma^{2}$, $\alpha_{1}$, and $\alpha_{2}$, of a level 22 DVG process, respectively. The dashed lines corresponds to Jan 25, May 11, Jun 14, and Nov 11.
@@
Figure 6: Estimated daily parameters of a level-2 DVG model for BTC log returns in 2022, with $90\%$ credible intervals obtained from the posterior distribution. The black line shows the posterior mean, while the blue dots indicate the posterior median.
@@
Figure 4 provides further insight by comparing characteristic functions and probability density functions for the selected high-volatility days. These plots illustrate the flexibility of the NBE approach in adapting to changing market conditions. The day-to-day variation in the characteristic and density functions underscores the nonstationary nature of financial data, and how the model dynamically adjusts to capture this complexity.
@@
Figure 5 offers a visual representation of the $90\%$ confidence intervals that surround the parameter estimates. Notably, the non-overlapping confidence intervals throughout the year indicate that the underlying process is indeed nonstationary – its parameters evolve meaningfully over time.
While the bootstrap procedure provides estimator uncertainty, the NBE framework also enables uncertainty quantification in a fully Bayesian manner. Figure 6 shows the estimates for the posterior mean, median, and illustrates the $90\%$ posterior-based credible intervals for the process variance of Bitcoin in 2022. Comparing Figures 5 and 6, we observe that the posterior credible intervals are generally wider than the bootstrap intervals, as they reflect full distributional uncertainty. This is particularly evident during periods of high volatility, where the posterior approach tends to be more conservative, while the bootstrap intervals remain relatively narrow in regions where the estimator is stable.
In particular, we define two cryptocurrencies as having different distributions on a given day if at least one of their corresponding confidence intervals does not overlap. Applying this criterion, we find that Bitcoin and XRP exhibit different distributions on 239 days in 2022, corresponding to roughly $65\%$ of the year. Similarly, Bitcoin and Ethereum differ on approximately $40\%$ of the days, and Ethereum and XRP differ on about $38\%$ of the days.
@@
Time [s] 45,552 656 5 (10.8 [hr]*)
Table 4: Time to perform daily parameter estimation for the year 2022 averaged across 3 different cryptocurrencies. * Training time for NBE. The reported 5 seconds refer to inference after training.
@@
In addition to accuracy and robustness, computational efficiency is a key consideration for practical deployment. Table 4 compares the estimation runtime of NBE against MELE and LSQ. NBE is approximately 9,000 times faster than MELE and 130 times faster than LSQ, making it an appealing option for real-time applications.
@@ 4.1 The Long-Term Evolution of Bitcoin
We also analyzed the long-term evolution of DVG process parameters fitted to BTC over a nine-year period spanning 2014 to 2022. The NBE approach shines in this application, with nearly a decade of data requiring only approximate 45 seconds to compute daily estimates for the entire period. Figure 7 displays these daily estimates for $\sigma^{2}$, $\alpha_{1}$, and $\alpha_{2}$ using the NBE framework. The parameters show clear signs of nonstationarity, with periods of relative stability punctuated by bursts of volatility. Since the variance estimates are difficult to interpret directly due to scale differences, we additionally visualize their evolution in log-scale in Figure 8.
@@ 5 Discussion
NBE is a flexible, amortized framework for parameter estimation, in which a neural network trained via simulation is used to approximate Bayes estimators, functionals that minimize the Bayes risk. This allows NBE to bypass the need for closed-form likelihoods, making it especially well-suited for complex models such as Lévy processes, where traditional methods are usually limited by computational cost.
Our work demonstrates that NBE not only provides computational efficiency and accuracy, but also comes with rigorous theoretical guarantees, including convergence in risk. Moreover, our simulation studies showed that NBE consistently achieves lower estimation error across all parameters compared to established techniques such as LSQ and MELE. Moreover, it does so with dramatically reduced computation times, achieving speedups of several orders of magnitude, making it a compelling option for large-scale or time-sensitive applications.
The choice between NBE and traditional methods such as LSQ or MELE depends, however, on the application context. If the goal is to analyze a single dataset and no further estimation is required, perhaps direct optimization via LSQ may be more practical, as it avoids the initial training phase of NBE. On the other hand, when estimation needs to be performed repeatedly, for example across multiple datasets, or large-scale simulation studies, the upfront training cost of NBE becomes negligible compared to the efficiency and accuracy gains at the inference stage. In such settings, NBE provides both faster and, as our results demonstrate, more accurate estimates. We refer the reader to Section 2 of the review paper Zammit-Mangion et al. [2025] which discusses this trade-off between training cost and inference efficiency.
We further demonstrated the utility of NBE on high-frequency cryptocurrency return data. The method effectively captured temporal changes in process parameters, responded sensitively to market events, and offered robust uncertainty quantification. In particular, we considered two complementary approaches for UQ: a nonparametric bootstrap procedure, which captures the sampling variability of the estimator, and a posterior-quantile approach, which directly approximates posterior quantiles and allows construction of credible intervals in a likelihood-free manner.
These findings highlight NBE’s capacity to adapt to the nonstationary nature of real-world financial data while delivering stable and interpretable estimates: we obtained parameter estimates for high-frequency 1-minute bitcoin data over 9 years in matter of seconds.
Nonetheless, certain limitations remain. Accurately estimating parameters associated with deep subordination, such as those in the level-2 DVG model, proved more difficult in low-signal regimes. Future work may focus on improving estimation for these challenging regimes, possibly by incorporating structured priors or attention mechanisms to enhance model sensitivity. In addition, future extensions could include more refined posterior approximations or alternative bootstrap strategies to enhance uncertainty assessment. Future directions may also include generalizing the method to multivariate Lévy processes, or exploring extrapolation beyond the prior distribution of the parameters.
@@ Acknowledgements
The authors thank the reviewers, associate editor and editor for constructive suggestions which improved the quality of the manuscript. This research was supported by National Science Foundation grant DMS-2310487.
@@ 6 Appendix
6.1 Proof of Theorem 2
We begin with the basic decomposition
$R(\hat{\theta}_{m,N}^{\star})-R^{\star}=\big(R(\hat{\theta}_{m}^{\star})-R^{\star}\big)+\big(R(\hat{\theta}_{m,N}^{\star})-R(\hat{\theta}_{m}^{\star})\big),$
where $R^{\star}=R(\hat{\theta}^{\star})$ is the true Bayes risk, $\hat{\theta}_{m}^{\star}$ denotes the population minimizer over the class $\mathcal{F}_{m}$, and $\hat{\theta}_{m,N}^{\star}$ its empirical counterpart based on $N$ samples.
Consider the first difference. By (A2) the Bayes estimator $\hat{\theta}^{\star}(\cdot)$ is permutation-invariant. The universality property of DeepSets (A3) ensures that for every $\varepsilon>0$ there exists $m$ and $\tilde{\theta}_{m}\in\mathcal{F}_{m}$ with
$\sup_{x}\|\tilde{\theta}_{m}(x)-\hat{\theta}^{\star}(x)\|\leq\varepsilon.$
Since $L$ is Lipschitz (A1), this implies for all $(\theta,x)$ that
$\big|L(\theta,\tilde{\theta}_{m}(x))-L(\theta,\hat{\theta}^{\star}(x))\big|\leq C_{L}\|\tilde{\theta}_{m}(x)-\hat{\theta}^{\star}(x)\|\leq C_{L}\varepsilon.$
Therefore,
$\begin{split}\left|R(\tilde{\theta}_{m}(x)))-R(\hat{\theta}^{\star}(x))\right|&=\big|\int_{\Theta}\int_{\mathcal{S}}\left[L(\theta,\tilde{\theta}_{m}(x))-L(\theta,\hat{\theta}^{\star}(x))\right]\,f(x\mid\theta)\,dx\,d\Omega(\theta).\big|\\ &\leq\int_{\Theta}\int_{\mathcal{S}}\left|L(\theta,\tilde{\theta}_{m}(x))-L(\theta,\hat{\theta}^{\star}(x))\right|\,f(x\mid\theta)\,dx\,d\Omega(\theta).\\ &\leq C_{L}\varepsilon\int_{\Theta}\int_{\mathcal{S}}\,f(x\mid\theta)\,dx\,d\Omega(\theta).\\ &=C_{L}\varepsilon.\end{split}$
Since $\hat{\theta}_{m}^{\star}$ is the minimizer of $R$ over $\mathcal{F}_{m}$,
$R(\hat{\theta}_{m}^{\star})\leq R(\tilde{\theta}_{m})\leq R(\hat{\theta}^{\star})+C_{L}\varepsilon.$
Thus $0\leq R(\hat{\theta}_{m}^{\star})-R^{\star}\leq C_{L}\varepsilon$, and since $\varepsilon>0$ was arbitrary, we conclude
$\lim_{m\to\infty}\big(R(\hat{\theta}_{m}^{\star})-R^{\star}\big)=0.$
For the second difference, fix $m$ and compare the empirical and population minimizers. Write
$R(\hat{\theta}_{m,N}^{\star})-R(\hat{\theta}_{m}^{\star})=\big(R(\hat{\theta}_{m,N}^{\star})-R_{N}(\hat{\theta}_{m,N}^{\star})\big)+\big(R_{N}(\hat{\theta}_{m,N}^{\star})-R_{N}(\hat{\theta}_{m}^{\star})\big)+\big(R_{N}(\hat{\theta}_{m}^{\star})-R(\hat{\theta}_{m}^{\star})\big).$
[…]
$0\leq R(\hat{\theta}_{m,N}^{\star})-R(\hat{\theta}_{m}^{\star})\leq 2\sup_{\hat{\theta}\in\mathcal{F}_{m}}|R_{N}(\hat{\theta})-R(\hat{\theta})|\xrightarrow{P}0,$
so the error vanishes as $N\to\infty$ for fixed $m$.
Finally, for any $\delta>0$, choose $m$ large enough that $R(\hat{\theta}_{m}^{\star})-R^{\star}\leq\delta$ by the approximation argument above. For this fixed $m$, choose $N$ large enough so that $R(\hat{\theta}_{m,N}^{\star})-R(\hat{\theta}_{m}^{\star})\leq\delta$ with high probability. Combining the two bounds yields
$R(\hat{\theta}_{m,N}^{\star})-R^{\star}\leq 2\delta,$
and since $\delta>0$ was arbitrary, we conclude that
$R(\hat{\theta}_{m,N}^{\star})\xrightarrow{P}R^{\star},$
for $(m,N)\to\infty$.
@@ 6.2 Simulation Studies
To train the estimator across a broad range of plausible values, we selected reasonably uninformative prior distributions for the parameters of each process, similar as the priors used in Section 3.3. These were chosen to reflect realistic yet flexible ranges for each model’s parameter space.
  • Compound Poisson process: $\lambda\sim U(0.1,1.3),\mu\sim U(-0.6,0.6),\sigma^{2}\sim U(10^{-3},0.3)$
  • Merton jump-diffusion model: $\mu\sim U(-0.8,0.8),\sigma^{2}\sim U(10^{-3},1)$, $\lambda\sim U(0.1,1.5)$, $\mu_{J}\sim U(-1.5,1.5),\sigma_{J}^{2}\sim U(0.1,1.7)$
  • Variance gamma process: $\gamma\sim U(-1.5,1.5),\sigma^{2}\sim U(10^{-4},2),\alpha\sim U(0.1,3)$
  • Deep variance gamma (level 2): $\sigma^{2}\sim U(10^{-6},3),\alpha_{1},\alpha_{2}\sim U(10^{-6},25)$
@@
6.2.1 Robustness to Priors
To assess the sensitivity of our estimates to the choice of prior, we conducted additional simulation experiments in which the prior distributions of each parameter was enlarged. As expected, the RMSE for the estimates tends to increase with wider priors, since the estimator now has bigger (in magnitude) plausible parameter values. To account for this effect and better evaluate consistency across prior specifications, we introduce two complementary metrics:
  • Normalized root-mean-square error (NRMSE):
    $\mathrm{NRMSE}(\hat{\theta})=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\hat{\theta}_{i}-\theta_{i})^{2}}}{\theta_{\mathrm{max}}-\theta_{\mathrm{min}}},$
    where $\theta_{\mathrm{max}}$ and $\theta_{\mathrm{min}}$ denote the upper and lower bounds of the prior for the parameter $\theta$. Normalizing by the prior length allows us to assess whether estimation quality remains proportional to the size of the parameter space.
  • Mean absolute percentage error (MAPE):
    $\mathrm{MAPE}(\hat{\theta})=\frac{1}{n}\sum_{i=1}^{n}\frac{|\hat{\theta}_{i}-\theta_{i}|}{|\theta_{i}|},$
    which provides a measure of average relative error for each parameter.
The importance of these metrics is that we argue the following. If the normalized RMSE and MAPE remain approximately stable under prior expansion, the method is robust and produces consistent estimates regardless of the specific prior choice.
For the robustness analysis, we considered three increasingly wide prior intervals for the parameters of the compound Poisson process. In the first specification (interval 1), we set $\lambda\sim U(0.1,1.3),\mu\sim U(-0.6,0.6),\sigma^{2}\sim U(10^{-3},0.3)$. To assess sensitivity, we then expanded the support of each parameter by a factor of five, yielding the second specification (interval 2), $\lambda\sim U(0.1,6.5),\mu\sim U(-3,3),\sigma^{2}\sim U(10^{-3},1.5)$, and finally by a factor of ten for the third specification (interval 3), $\lambda\sim U(0.1,13),\mu\sim U(-6,6),\sigma^{2}\sim U(10^{-3},3)$.
Similarly, we conducted the experiment with the DVG process of level 2. In the first specification (interval 1), we use the same bounds as in the applications section, $\sigma^{2}\sim U(10^{-6},3),\alpha_{1},\alpha_{2}\sim U(10^{-6},25).$ Then, we expanded the support of each parameter by a factor of five, yielding the second specification (interval 2), $\sigma^{2}\sim U(10^{-6},6),\alpha_{1},\alpha_{2}\sim U(10^{-6},50)$.
As reported in Tables 5 and 6, while the absolute RMSE increases with wider priors, the NRMSE and MAPE for NBE remain relatively constant, suggesting that the neural Bayes estimator maintains reliable performance even under prior expansion. This provides empirical evidence of robustness to prior specification. Moreover, the training times across the different settings were essentially comparable and thus did not represent a meaningful change in computational cost.
Parameter/metric Interval 1 Interval 2 Interval 3
$\lambda$ NBE LSQ NBE LSQ NBE LSQ
RMSE 0.051 0.058 0.33 1.38 1.11 3.96
Bias 0.0005 -0.004 -0.0042 0.52 -0.012 1.54
NRMSE 0.051 0.058 0.053 0.22 0.087 0.31
MAPE 0.052 0.057 0.063 0.26 0.1 0.66
$\mu$ NBE LSQ NBE LSQ NBE LSQ
RMSE 0.028 0.23 0.13 2.47 0.31 4.71
Bias -0.005 -0.02 -0.045 -0.05 -0.06 -0.05
NRMSE 0.023 0.19 0.021 0.41 0.026 0.39
MAPE 0.22 0.46 0.16 1.22 0.18 3.44
$\sigma^{2}$ NBE LSQ NBE LSQ NBE LSQ
RMSE 0.022 0.13 0.20 0.37 0.62 1.23
Bias 7.6e-5 0.12 0.038 0.0019 -0.06 -0.56
NRMSE 0.092 0.55 0.14 0.26 0.21 0.42
MAPE 0.11 0.97 0.39 0.71 0.72 0.87
Table 5: Comparison of RMSE, Bias, NRMSE, and MAE (averaged over 1000 simulated datasets from the prior) for the parameters of the compound Poisson process across three intervals . The results are averaged and obtained using $nt=500$, $K=5000$, and $J=10$.
Parameter/metric Interval 1 Interval 2
$\sigma^{2}$ NBE LSQ NBE LSQ
RMSE 0.051 0.08 0.20 1.25
Bias 0.0024 0.0031 -0.011 0.008
NRMSE 0.017 0.02 0.014 0.08
MAPE 0.025 0.02 0.023 0.073
$\alpha_{1}$ NBE LSQ NBE LSQ
RMSE 6.72 9.63 35.34 53.8
Bias -2.17 -0.66 -11.57 6.69
NRMSE 0.27 0.39 0.28 0.43
MAPE 0.75 1.78 1.00 2.35
$\alpha_{2}$ NBE LSQ NBE LSQ
RMSE 6.72 10.4 36.36 53.12
Bias -2.54 2.3 -12.68 9.8
NRMSE 0.27 0.42 0.29 0.42
MAPE 0.66 1.00 0.99 2.06
Table 6: Comparison of RMSE, Bias, NRMSE, and MAE (averaged over 1000 simulated datasets from the prior) for the parameters of a DVG process of level 2 across two intervals. The results are averaged and obtained using $nt=500$, $K=5000$, and $J=10$.
@@ 6.2.2 Experiments on $n_{t}$
We began our investigation by fixing K=5, [^0] K=5{,}000 and $J=10$, and exploring the impact of $n_{t}$—the number of observations per dataset—on both training time and estimation accuracy. Specifically, we evaluated how varying $n_{t}$ affects the root mean squared error (RMSE) of the estimated parameters and the overall training time of the neural network.
To test robustness, we repeated this analysis across the three Lévy process models discussed in Section 3.2. For each model and each value of $n_{t}$, the estimator was trained on $K\cdot J$ datasets and then evaluated on an independent test set consisting of $1,000$ simulated datasets.
To visualize the results, we plotted RMSE against $n_{t}$ for each parameter in each model in Figure 9. These plots help identify the point at which increasing the input length yields little or no additional improvement in accuracy. Across all models, we observe that RMSE steadily improves as $n_{t}$ increases, with diminishing returns beyond approximately 3,0003{,}0005,0005{,}000 observations. The figure also presents results for two different models—the compound Poisson and variance gamma processes—demonstrating similar trends. In fact, for the variance gamma process (bottom row), the point of diminishing returns appears even earlier, around 1,5001{,}500 observations. Figure 10 also shows how the training time increases as a function of $n_{t}$.
Given these findings, we selected $n_{t}=1{,}440$ in our final experiments. This value corresponds to the number of one-minute log returns in a single trading day (24 hours), making it a natural and comparable choice for our cryptocurrency data application. While it lies slightly below the saturation threshold, it strikes a practical balance between empirical performance and application-specific relevance.
@@
Figure 10: Time in seconds to train the NBE for the Compound Poisson (left) and Variance Gamma (right) processes as a function of input length $n_{t}$.